# Leveraging Conditional Mutual Information to Improve Large Language Model Fine-Tuning For Classification

### v1

@@ II-B LLMs as High-Order Markov Source Models
In the above, we omit the operation details of each transformer block ${\cal T}_{l}$, which involve matrix and other computation with its own parameters $\Theta_{l}$. A very important factor is that all transform blocks ${\cal T}_{l}$ are causal [^1] This causality does not apply to transformer encoder-based LLMs such as BERT [4] which are mainly used for downstream tasks.. This, in turn, ensures that the probability distribution computed in (3) indeed can be interpreted as a conditional distribution given the present token $t_{i}$ and past tokens $t_{1}\cdots t_{i-1}$, but not the future tokens.
Since at the beginning, the first token is often a special token, denoted as $[Start]$ and acting as an anchor point for initializing a model, the LLM architecture described above is equivalent, mathematically, to a family of $k$-th order homogeneous Markov source models $\{P(\cdot|t_{1},\dots,t_{i};(W_{e},W_{p},\{\Theta_{l}\}_{l=1}^{L})):1\leq i\leq k\}$ parameterized by $\Theta=(W_{e},W_{p},\{\Theta_{l}\}_{l=1}^{L})$. A specific LLM is then determined by a training process which tries to minimize over $\Theta$,
${\cal L}_{1}({\cal U})=\sum_{(u_{1},\dots,u_{k},u_{k+1})\in{\cal U}}-\sum_{i=2}^{k+1}\log P(u_{i}|u_{1},\dots,u_{i-1};\Theta)$ (4)
@@ IV Experiments
In this section, we present the experiment results on LLM fine-tuning for classification. While the proposed CMI methods with new definitions of CMI could also be applied during the pretraining phase of LLM development—where the number of centroids would correspond to the model’s vocabulary size of over 30,000—this stage entails significantly higher computational complexity. Due to resource constraints, our experiments instead focus on the fine-tuning stage, which offers a more practical approach to evaluating the effectiveness of CMI in LLM-based classification.
@@ IV-A Minimizing CMI for Model Optimization
To evaluate the role of CMI minimization in optimizing LLMs for standalone performance in classification tasks, we conducted experiments on the BERT-base model [4] for all classification tasks from the GLUE benchmark [23]. Each model was initialized with its pretrained weights and was fine-tuned by minimizing its CMI as explained in Section III-C. This approach will encourage the model to focus on task-relevant features while reducing noise introduced by unnecessary contextual dependencies.

### v2

@@ II-B LLMs as High-Order Markov Source Models
In the above, we omit the operation details of each transformer block ${\cal T}_{l}$, which involve matrix and other computation with its own parameters $\Theta_{l}$. An important factor is that all transformer blocks ${\cal T}_{l}$ are causal, unlike in encoder-based models such as BERT [4], which are primarily designed for downstream tasks. This, in turn, ensures that the probability distribution computed in (3) indeed can be interpreted as a conditional distribution given the present token $t_{i}$ and past tokens $t_{1}\cdots t_{i-1}$, but not the future tokens.
Since at the beginning, the first token is often a special token, denoted as $[Start]$ and acting as an anchor point for initializing a model, the LLM architecture described above is equivalent, mathematically, to a family of $k$-th order homogeneous Markov source models $\{P(\cdot|t_{1},\dots,t_{i};(W_{e},W_{p},\{\Theta_{l}\}_{l=1}^{L})):1\leq i\leq k\}$ parameterized by $\Theta=(W_{e},W_{p},\{\Theta_{l}\}_{l=1}^{L})$. A specific LLM is then determined by a training process (called cross-entropy) which tries to minimize over $\Theta$,
${\cal L}_{1}({\cal U})=\sum_{(u_{1},\dots,u_{k+1})\in{\cal U}}-\sum_{i=2}^{k+1}\log P(u_{i}|u_{1},\dots,u_{i-1};\Theta)$ (4)
@@ IV Experiments
In this section, we present the experiment results on LLM fine-tuning for classification. While the proposed CMI methods with new definitions of CMI could also be applied during the pretraining phase of LLM development (where the number of centroids would correspond to the model’s vocabulary size of over 30,000), this stage entails significantly higher computational complexity. Due to resource constraints, our experiments instead focus on the fine-tuning stage, which offers a more practical approach to evaluating the effectiveness of CMI in LLM-based classification.
@@ IV-A Minimizing CMI for Model Optimization
To evaluate the role of CMI minimization in optimizing LLMs for standalone performance in classification tasks, we conducted experiments on the BERT-base model [4] for all classification tasks from the GLUE benchmark [23], a suite of language understanding tasks. Each model was initialized with its pretrained weights and was fine-tuned by minimizing its CMI as explained in Section III-C. This approach will encourage the model to focus on task-relevant features while reducing noise introduced by unnecessary contextual dependencies.